The data from a calcium binding experiment are provided. Calcium is the log of free calcium concentration and ProteinProp is the proportion of the protein bound the calcium.

Calcium ProteinProp
-10.14539036 0.145164202
-9.97798426 0.223711498
-9.351249787 0.219828821
-9.101000729 0.334269431
-9.01376587 0.378526213
-8.940436582 0.409369052
-8.578231599 0.507444972
-8.370182804 0.57164134
-8.289036881 0.642186965
-7.959793372 0.807279963
-7.592269272 0.930025173
-7.238448011 0.901409641
-7.038626372 0.950327606
-6.330776126 0.957320456
-6.167235695 0.985104489
-5.556893543 0.969407033
-5.321208566 0.999285187
-4.813608784 1
-10.14539036 0.188284071
-9.97798426 0.226840832
-9.351249787 0.299825094
-9.101000729 0.351716328
-9.01376587 0.413916136
-8.940436582 0.437475476
-8.578231599 0.526377071
-8.370182804 0.619739975
-8.289036881 0.670996494
-7.959793372 0.844443513
-7.592269272 0.929812237
-7.238448011 0.979803174
-7.038626372 0.974212892
-6.330776126 0.974230866
-6.167235695 0.987524724
-5.556893543 0.998230049
-5.321208566 1
-4.813608784 0.995714469
-10.72193267 0.264766386
-10.44575319 0.336968113
-9.689731633 0.401104
-9.047837426 0.397172676
-8.791558644 0.535642225
-8.448916135 0.648687709
-8.08820341 0.668027433
-7.851397345 0.805547503
-7.658565475 0.858684464
-7.482276405 0.879804722
-7.306448914 1
-7.115544504 0.977186242
-6.884056823 0.965169616
-6.539854183 0.964522047
-5.86518563 0.98589626

a. Fit a quadratic regression model for predicting ProteinProp from Calcium. Write down the fitted regression equation.of

b. Add the quadratic curve to a scatterplot of ProteinProp versus Calcium

c. Are the conditions for inference reasonably satisfied for this model?

d. Is the parameter for the quadratic term significantly fifferent from zero? Justify.

e. Identify the coefficient of multiple determination and interpret this value
Answer #1

a.The regression equation

Y(ProteinPrep)= 1.429 - 0.011*(Calcium^2)

c.the conditions for inference reasonably satisfied for this model

d.the pvalue is approximately 0 which means the parameter is not significantly 0

e.Coeff of multiple determination = R-Sq=0.8741

which means 87.41% of total variation of proteinPrep is explained by calcium

Please Give A Thumbs Up